OpenStudy (hockeychick23):

Pre-calc help please? During a research experiment, it was found that the number of bacteria in a culture grew at a rate proportional to its size. At 5:00 AM there were 5,000 bacteria present in the culture. At noon, the number of bacteria grew to 5,600. How many bacteria will there be at midnight?

OpenStudy (hockeychick23):

\[5000e^7=5600 --> 7k= \ln(5600/5000)> t(14)=5000e^.01618981\] @Photon336 this is as far as I got can you help me please?

OpenStudy (hockeychick23):

@ILovePuppiesLol can you help me with this please?

OpenStudy (hockeychick23):

@jim_thompson5910 can you help me with this please?

jimthompson5910 (jim_thompson5910):

so they gave you a formula or did they want you to derive one out?

OpenStudy (hockeychick23):

they wanted me to derive one

jimthompson5910 (jim_thompson5910):

I'm guessing you haven't learned about derivatives?

OpenStudy (hockeychick23):

not really no

jimthompson5910 (jim_thompson5910):

hmm let me think

jimthompson5910 (jim_thompson5910):

did your teacher give you a template to work with?

OpenStudy (hockeychick23):

A= P(1+r/n)^n/t and A= Pe^rt

jimthompson5910 (jim_thompson5910):

ok it sounds like we'll be using A= Pe^(rt)

jimthompson5910 (jim_thompson5910):

A= Pe^(rt) A= Pe^(r*0) ... plug in t = 0 for time 0 5000 = P*e^(r*0) .... plug in the population at time 0 solve for P to get P = 5000 agreed so far?

OpenStudy (hockeychick23):

yep

jimthompson5910 (jim_thompson5910):

ok so A = 5000*e^(r*t) is the equation so far

OpenStudy (hockeychick23):

yea

jimthompson5910 (jim_thompson5910):

now plug in A = 5600 and t = 7 (noon is seven hours after 5 AM) A = 5000*e^(r*t) 5600 = 5000*e^(r*7) then solve for r

jimthompson5910 (jim_thompson5910):

it looks like you were doing that at the top now that I look it over more thoroughly

OpenStudy (hockeychick23):

oh ok, well i got .01618981 for r

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

the equation is now A = 5000*e^(0.01618981*t)

jimthompson5910 (jim_thompson5910):

plug in t = 19 (add 12 more hrs to t = 7 to get t = 19 to get to midnight) A = 5000*e^(0.01618981*t) A = 5000*e^(0.01618981*19) A = ??

OpenStudy (hockeychick23):

i got 6800.827

jimthompson5910 (jim_thompson5910):

so roughly 6801 bacteria

OpenStudy (hockeychick23):

ok thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome